A small aperture is illuminated with a parallel beam of λ = 628nm . The emergent beam has an angular divergence of 2∘ . The size of the aperture is.
1
Expert's answer
2019-09-03T10:25:42-0400

Relation between an angular divergence(in radians), wavelength and size of the aperture:


"\alpha = \frac{\lambda}{d}"

Therefore,


"d = \frac{\lambda}{\alpha} = \frac{628\text{nm}}{2\cdot \pi / 180} = 18\mu m"